OpenStudy (anonymous):

For the hyperboloid whose equation is x=4z^2-y^2+2y derive the scalar eqn for the plane that is tagent at pt (4,2,-1)?

OpenStudy (anonymous):

Okay is this just as simple as I think it is using the tangent equation z=f(a,b,c) + f_x(x-a) + f_y(y-b)+f_z(z-c)

OpenStudy (anonymous):

any ideas sirm3d?

OpenStudy (sirm3d):

first step is to rewrite the equation of the surface to F(x,y,z)=0, then evaluate the first partial derivatives at the specified point.

OpenStudy (anonymous):

Okay I got f_x = -1, f_y=-2y-2 = -2 and f_z=8z = -8

OpenStudy (sirm3d):

compute these values \[\large F_x(4,2,-1),F_y(4,2,-1), F_z(4,2,-1)\] for the function \[\large F(x,y,z)=x-4z^2+y^2-2y\]. can you give me the result of your computation?

OpenStudy (anonymous):

do I use the tangent equation now?

OpenStudy (sirm3d):

uhm, you mean \[\large f_y=-2y+2=-2\]

OpenStudy (anonymous):

yeah sorry minor typo

OpenStudy (sirm3d):

if a plane has the vector representation <a, b, c>, the scalar form of the plane through (x0, y0, z0) is \[\large a(x-x_0)+b(y-y_0) + c(z-z_0) = 0\]

OpenStudy (anonymous):

Ok so the vector representation in this case is that the f_x, f_y, f_z?

OpenStudy (anonymous):

To give -(x-4) -2(y-2) -8(z+1)?

OpenStudy (sirm3d):

right

OpenStudy (sirm3d):

equated to zero, of course.

OpenStudy (anonymous):

Do you know why though? Is it because since we're taking the derivative for each case

OpenStudy (anonymous):

actually never mind I don't know why haha

OpenStudy (anonymous):

vector representation is just the gradient?

OpenStudy (anonymous):

this looks very similar to the tangent plane equation

OpenStudy (sirm3d):

the gradient (at a point) of the function is the vector representation of the tangent plane (at that point).

OpenStudy (anonymous):

Oh gotcha, okay awesome explanation thanks a lot for your help! Very much appreciated